OpenStudy (anonymous):

Anyone fancy earning themselves a medal? quick question... what is ln7x=2ln9x

OpenStudy (mr.math):

I don't! Can I still answer?

OpenStudy (anonymous):

yes, go ahead (:

OpenStudy (mr.math):

Is it \(\ln(7x)=2\ln(9x)?\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

please give your answer to 3 d.p

OpenStudy (mr.math):

Okay. \[\ln(7x)-2\ln(9x)=0 \implies \ln\frac{7x}{(9x)^2}=0 \implies \frac{7}{81x}=1\] \[\implies 81x=7 \implies x=\frac{7}{81}.\]

OpenStudy (mr.math):

\[x=\frac{7}{81}\approx 0.086\]

OpenStudy (anonymous):

thanks , shame you didn't want a medal!

OpenStudy (mr.math):

Lol, you can give it though :P